An advertising agency conducted a survey and found that the number of units sold, N, is related to the amount a spent on advertising (in dollars) by the following formula:
N = 1,500 + 100 ln a (a ≥ 1)
How many units are sold after spending $1,000? (Round your answer to the nearest whole number.)
units

Answer by stanbon(75887)

An advertising agency conducted a survey and found that the number of units sold, N, is related to the amount a spent on advertising (in dollars) by the following formula:
N(a) = 1,500 + 100 ln a (a ≥ 1)
How many units are sold after spending $1,000? (Round your answer to the nearest whole number.)
--------------
N(1000) = 1,500 + 100*ln(1000) = 1500 + 690.78 = 2191
===============
Cheers,
Stan H.
===============